# In defense of temporal Tsirelson bound

### v1

@@ Abstract
In a recent paper, Chatterjee et al. [Phys. Rev. Lett 135, 220202 (2025)] analyze and experimentally implement a specific unitary evolution of a simple quantum system. The authors refer to this type of dynamics as a “superposition of unitary time evolutions.” They claim that such an evolution enables a violation of the temporal Tsirelson bound in the Leggett–Garg scenario, a claim that is supported by their experimental results. In this work, we show that the proposed evolution can be understood within a more conventional framework, without invoking a superposition of evolutions. Furthermore, we demonstrate that the apparent violation of the bound arises because the measured quantities are not consistent with the assumptions of the Leggett–Garg scenario. This is a slightly extended version of the Comment submitted for publication in Phys. Rev. Lett.
In a recent paper, Chatterjee et al.Chatterjee (2025) define operator $\mathcal{U}(t_{f},t_{0})$ (see Eqs.(1) and (2) in Chatterjee (2025)) acting on 2-dimensional quantum systems. Here, we will use slightly different notation $U_{p}(t)=\mathcal{U}(t_{f},t_{0})$, where $p=(\phi,\alpha)$, $t=\omega(t_{f}-t_{0})/2$ using parameters ($\phi,\alpha,\omega,t_{f},t_{0}$) introduced in Chatterjee (2025). Our notation emphasizes the fact that $\mathcal{U}(t_{f},t_{0})$ depends only on the difference $t_{f}-t_{0}$. Chatterjee et al.Chatterjee (2025) construct $U_{p}(t)$ in the following way. First $\sigma_{\phi}$ is defined as a linear combination of Pauli operators
@@
After observation that $W_{p}^{\dagger}(t)W_{p}(t)$ is proportional to identity $W_{p}^{\dagger}(t)W_{p}(t)=n_{p}^{2}(t)I$, where
@@ I Leggett-Garg scenario
Let us now consider the three-term Leggett–Garg scenario. It consists of three distinct experiments (see Fig. 3). All three share the following elements: three fixed times $t_{1}<t_{2}<t_{3}$, the same initial state prepared at $t=t_{1}$, and the same unitary evolution operators $V_{12}$ and $V_{23}$. In each experiment, the unitary evolution is interrupted at two distinct times by measurements of the same dichotomic observable $Q$. This allows one to estimate correlators $C_{12}$, $C_{23}$, and $C_{13}$, where $C{jk}=\left<Q(t_{j})Q(t_{k})\right$>. One can then compute the Leggett–Garg parameter $K_{3}=C_{12}+C_{23}-C_{13}$. The temporal Tsirelson bound (TTB) is given by the inequality $K_{3}\leq\frac{3}{2}$. Let us now concentrate on the specific case presented in Chatterjee (2025). Their initial state was a maximally mixed state $\frac{I}{2}$ and observable $Q=\sigma_{z}$. This allows us to write $C_{jk}=\frac{1}{2}tr\left[\sigma_{z}V_{jk}\ \sigma_{z}V_{jk}^{\dagger}\right]$. A crucial requirement of the Leggett–Garg scenario is that the unitary evolution be identical in all three experiments. So, it must be
@@ II Modifications to the Experimental Protocol
In both Scenarios A and B, the composition condition (Eq. 18) is strictly satisfied. Consequently, we find no violation of the Temporal Tsirelson Bound (TTB) in either case.

### v2

@@ Abstract
In a recent paper, Chatterjee et al. [Phys. Rev. Lett 135, 220202 (2025)] analyze and experimentally implement a specific unitary evolution of a simple quantum system. The authors refer to this type of dynamics as a “superposition of unitary time evolutions.” They claim that such an evolution enables a violation of the temporal Tsirelson bound in the Leggett–Garg scenario, a claim that is supported by their experimental results. In this work, we show that the proposed evolution can be understood within a more conventional framework, without invoking a superposition of evolutions. Furthermore, we demonstrate that the apparent violation of the bound arises because the measured quantities are not consistent with the assumptions of the Leggett–Garg scenario.
It is well established that a classical description of the physical world is incomplete. One of the most striking examples of this incompleteness is the limit imposed on correlations between measurement outcomes. Bell Bell (1964) and Leggett-Garg Leggett and Garg (1985) famously formulated inequalities that bound these correlations in space and time, respectively. Quantum mechanics, however, violates these bounds (for review see Brunner et al. (2014); Emary et al. (2014)), standing in stark contrast to the predictions of classical theory. These violations are not limitless and their bounds are usually called the Tsirelson bound and temporal Tsirelson bound for Bell and Leggett-Garg scenarios respectively.
Although temporal Tsirelson bound is well established within scientific community, in a recent paper, Chatterjee et al.Chatterjee (2025) claim to experimentally violate it. Further the Authors claim that this violation is based on proposed new kind of quantum evolution. In this note we show to the contrary that not only the presented evolution is nothing more than a simple unitary transformation but also that the violation of temporal Tsirelson bound is only apparent.
To start with let us define operator $\mathcal{U}(t_{f},t_{0})$ as Eqs.(1) and (2) in Chatterjee (2025) acting on 2-dimensional quantum systems. However, we will use slightly different notation $U_{p}(t)=\mathcal{U}(t_{f},t_{0})$, where $p=(\phi,\alpha)$, $t=\omega(t_{f}-t_{0})/2$ using parameters ($\phi,\alpha,\omega,t_{f},t_{0}$) introduced in Chatterjee (2025). Our notation emphasizes the fact that $\mathcal{U}(t_{f},t_{0})$ depends only on the difference $t_{f}-t_{0}$. Chatterjee et al.Chatterjee (2025) construct $U_{p}(t)$ in the following way. First $\sigma_{\phi}$ is defined as a linear combination of Pauli operators
@@
After observation that $W_{p}^{+}(t)W_{p}(t)$ is proportional to identity $W_{p}^{+}(t)W_{p}(t)=n_{p}^{2}(t)I$, where
@@ I Leggett-Garg scenario
Let us now consider the three-term Leggett–Garg scenario. It consists of three distinct experiments (see Fig. 3). All three share the following elements: three fixed times $t_{1}<t_{2}<t_{3}$, the same initial state prepared at $t=t_{1}$, and the same unitary evolution operators $V_{12}$ and $V_{23}$. In each experiment, the unitary evolution is interrupted at two distinct times by measurements of the same dichotomic observable $Q$. This allows one to estimate correlators $C_{12}$, $C_{23}$, and $C_{13}$, where $C{jk}=\left<Q(t_{j})Q(t_{k})\right$>. One can then compute the Leggett–Garg parameter $K_{3}=C_{12}+C_{23}-C_{13}$. The temporal Tsirelson bound is given by the inequality $K_{3}\leq\frac{3}{2}$ Jordan et al. (2006); Kofler and Brukner (2008); Barbieri (2009); Budroni et al. (2013). Let us now concentrate on the specific case presented in Chatterjee (2025). Their initial state was a maximally mixed state $\frac{I}{2}$ and observable $Q=\sigma_{z}$. This allows us to write $C_{jk}=\frac{1}{2}tr\left[\sigma_{z}V_{jk}\ \sigma_{z}V_{jk}^{\dagger}\right]$. A crucial requirement of the Leggett–Garg scenario is that the unitary evolution be identical in all three experiments. So, it must be
@@ II Modifications to the Experimental Protocol
In both Scenarios A and B, the composition condition (Eq. 18) is strictly satisfied. Consequently, we find no violation of the Temporal Tsirelson Bound in either case.